OpenStudy (anonymous):

Basic calc question. I am given the derivative and the value of y at a given x value, can i find this?

OpenStudy (anonymous):

I know g(5)=-3 and g(5)'=6. I am asked to find the derivative of f(5). I am told f(x)=[g(x)]^3 . Is this capable of being solved, and if so how.

OpenStudy (rational):

differentiate using chain rule

OpenStudy (rational):

f(x) = [g(x)]^3 f'(x) = ?

OpenStudy (anonymous):

I guessed that but what is the derivative of a floating exponent. Im sorry if im typing slowly for some reason this site is killing my computer.

OpenStudy (anonymous):

6*3*-3^2?

OpenStudy (rational):

use this \[\frac{d}{dx}x^n ~=~n*x^{n-1}\]

OpenStudy (anonymous):

Yeah thats like the first rule of differntation you learn, but does it still appy without a varible x?

OpenStudy (rational):

variables are just dummy

OpenStudy (rational):

\[\large \frac{d}{d\clubsuit} \clubsuit^n ~= ~n*\clubsuit^{n-1} \]

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

S0 6*3*(-3)^2 would fit that or not?

OpenStudy (rational):

\[ f(x) = [g(x)]^3\] \[f'(x) = 3*[g(x)]^{3-1} * \color{red}{g'(x)}\]

OpenStudy (rational):

plugin \(x=5\)

OpenStudy (rational):

6*3*(-3)^2 looks good!